OpenStudy (anonymous):

is this correct? solve the equation in the real number system x^4-2x^3+8x^2-32x-126=0 answer: -16 x^5+x^4-32 x-126 = 0

OpenStudy (anonymous):

or is there no solution

OpenStudy (anonymous):

how the heck are you supposed to solve the zeros of a fourth degree polynomial without technology? http://www.wolframalpha.com/input/?i=+x^4-2x^3%2B8x^2-32x-126%3D0+

OpenStudy (anonymous):

that shows so many answers how do i know the right one

OpenStudy (anonymous):

math is extremely hard for me sorry

OpenStudy (anonymous):

there are only two real solutions listed

OpenStudy (anonymous):

i understand frankly it is almost impossible to solve this you need a computer

OpenStudy (anonymous):

wolfram lists two real solutions and two complex solutions how were you supposed to solve this?

OpenStudy (anonymous):

it says my answer should be x= or no solution yeah i wish i knew all of this

OpenStudy (anonymous):

are you sure there is not a typo in your question? the answers are very ugly

OpenStudy (anonymous):

i have about 15 questions i dont understand and due tonight :(

OpenStudy (anonymous):

let me double check

OpenStudy (anonymous):

can you link or take a screen shot?

OpenStudy (anonymous):

it wont let me for some reason i cant even copy and paste here is the question again it is possible i messed it up. solve the equation in real number system. x^4-2x^3+8x^2-32x-128=0

OpenStudy (anonymous):

that's more like it, \(-2\) and \(4\)

OpenStudy (anonymous):

how many more?

OpenStudy (anonymous):

can i just give you a hundred medals lol. how many more ?

OpenStudy (anonymous):

lets knock a few out quick

OpenStudy (anonymous):

in the spirit of "get it over with, i'll understand it some time later"

OpenStudy (anonymous):

questions is that what your talkin bout cause i have to open new ones for medals for you!

OpenStudy (anonymous):

i get no frequent flyer miles for medals, but if you want to start a new post, fine

OpenStudy (anonymous):

otherwise just ask here, when i get tired i'll let you know

OpenStudy (anonymous):

solve the inequality algebraically 4x-4>-3x^2. your awesome thank you

OpenStudy (anonymous):

start with \[3x^2+4x-4>0\]

OpenStudy (anonymous):

then factor as \[(3x-2)(x+2)>0\]

OpenStudy (anonymous):

zeros are at \(-2\) and \(\frac{2}{3}\) and since this is a parabola that faces up , it is positive outside the zeros, namely if \(x<-2\) or \(x>\frac{2}{3}\)

OpenStudy (anonymous):

if you need interval notation it is \[(-\infty,-2)\cup (\frac{2}{3},\infty)\]

OpenStudy (anonymous):

12 to go

OpenStudy (anonymous):

your so quick ;)

OpenStudy (anonymous):

half quick

OpenStudy (anonymous):

million times quicker than me....solve the fallowing inequality 56x-15>56/x

OpenStudy (anonymous):

i wish i could repay you cause this is the fail or pass moments

OpenStudy (anonymous):

it is really \[56x-15>\frac{56}{x}\]??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

lets try to make sure you pass is that the question for real?

OpenStudy (anonymous):

let m check again

OpenStudy (anonymous):

this is a pain, but we can do it. you need the whole method (like if you have to take an exam?) or just the answer?

OpenStudy (anonymous):

there is more to it sorry

OpenStudy (anonymous):

\[-\frac{7}{8}<x<0\] or \[x>\frac{8}{7}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the point of x=0 included in the solution? are there other finite points of the interval in the solution set? what is the solution set or is there no solution?

OpenStudy (anonymous):

yes, 0 is excluded because you have a zero in the denominator

OpenStudy (anonymous):

0 is included?

OpenStudy (anonymous):

EX cluded

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

solution is the one i wrote above

OpenStudy (anonymous):

\[-\frac{7}{8}<x<0\] or \[x>\frac{8}{7}\]

OpenStudy (anonymous):

this are there other finite points of the interval in the solution set? makes no sense to me , but i assume the answer is "no"

OpenStudy (anonymous):

11 to go

OpenStudy (anonymous):

use the remainder theorem to find the remainder when f(x) is divided by x-1/5. then use the factor theorem to determine whether x-1/5 is a factor of f(x). f(x)=5x^4-x^3-5x+1

OpenStudy (anonymous):

christ you math teacher must hate you

OpenStudy (anonymous):

im sure she does have never met her too lol

OpenStudy (anonymous):

do you need to show your work, or just answer the question? the answer is "the remainder is 0, because \(f(\frac{1}{5})=0\)"

OpenStudy (anonymous):

and therefore "yes, \(x-\frac{1}{5}\) is a factor"

OpenStudy (anonymous):

show work :(

OpenStudy (anonymous):

ok then compute \[f(\frac{1}{5})=5(\frac{1}{5})^4-(\frac{1}{5})^3-5(\frac{1}{5})+1=0\]

OpenStudy (anonymous):

it is actually not that hard you get \[\frac{1}{5^3}-\frac{1}{5^3}-1+1=0\]

OpenStudy (anonymous):

wow this is crazy

OpenStudy (anonymous):

therefore the remainder is zero and \(x-\frac{1}{5}\) is a factor finished

OpenStudy (anonymous):

10 to go

OpenStudy (anonymous):

when is this due?

OpenStudy (anonymous):

lets see how many we can finish in half an hour i bet we can do them all

OpenStudy (anonymous):

due tonight have to decide if im gonna pass or fail.im down to finish in half hour ;).. use the rational zeros theorem to list the potential rational zeros of the polynomial function do not attempt to find zeros f(x)= -25x^9-x^8+x+5

OpenStudy (anonymous):

numerator has to divide 5, denominator has to divide 25 there are not that many \[\pm1,\pm5,\pm\frac{1}{5},\pm\frac{1}{25}\]

OpenStudy (anonymous):

don't forget the \(\pm\)

OpenStudy (anonymous):

9 more, lets knock em out so you don't have to retake it

OpenStudy (anonymous):

use thw intermediate value theorem to show that the polynomial function has a zero in the given interval. f(x)=x^5-x^4+5x^3-4x^2-17x+8 ; [1.5, 1.8] find value s of 1.5 and 1.8

OpenStudy (anonymous):

\[f(1.5)=-7.09375\]

OpenStudy (anonymous):

\[f(1.8)=1.99808\]

OpenStudy (anonymous):

since one is positive, and the other is negative, there must be a number between 1.5 and 1.8 for which \(f\) is zero 8 more

OpenStudy (anonymous):

information is given about a polynomial f(x) whose coefficients are real numbers. find remaining zeros of f. degree3 ; zeros:1,-5-i

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

wow missed it somehow

OpenStudy (anonymous):

zeros are \[3, -5-i, -5+i\]

OpenStudy (anonymous):

polynomial is \[(x-3)(x-(-5+i))(x-(-5-i))\] \[=x^3+7 x^2-4 x-78\]

OpenStudy (anonymous):

7 more

OpenStudy (anonymous):

this one is a hard one

OpenStudy (anonymous):

did i loose ya?

OpenStudy (anonymous):

oh wow my computer did something weird so sorry

OpenStudy (anonymous):

actually i only have 1 left

OpenStudy (anonymous):

similar to the last :)

OpenStudy (anonymous):

I sat watching waiting for a response and i had to refresh everything after a long time

OpenStudy (anonymous):

if your still up to it here is the last one i will ask you for :) form a polynomial f(x) with real coefficients having the given degree and zeros. degree; 5 zeros: -4 ; i ; -2+i